# Phase-Imaging Ion-Cyclotron-Resonance Mass Spectrometry with the Canadian Penning Trap at CARIBU

### v4

@@ 1 Introduction
The Canadian Penning Trap (CPT) [4, 5] is located at Argonne National Laboratory (ANL) where it has conducted precision mass measurements of neutron-rich nuclides from the CAlifornium Rare Isotope Breeder Upgrade (CARIBU) [6] at the Argonne Tandem Linac Accelerator System (ATLAS) facility, using the Phase Imaging Ion-Cyclotron-Resonance (PI-ICR) [7, 8] technique. This paper discusses the PI-ICR operations, and some new systematic studies which have enabled the CPT to achieve ppb-level precision in mass measurements.
@@ 3.4 Penning trap
. Electrodes Voltages [V] Capture Trap Eject Bottom correction tube --17.5 --0.9413 --0.9413 Bottom end cap --13.5 --3.0836 --3.0836 Bottom correction ring --8.9622 --8.9622 --8.9622 Ring --12.7206 --12.7206 --12.7206 Top correction ring --8.9622 --8.9622 --8.9622 Top end cap --3.0836 --3.0836 --13.5 Top correction tube --0.9413 --0.9413 --17.5
@@ 4.1 PI-ICR at the CPT
PI-ICR relies on determining a nuclide’s $\nu_{c}$ via a “sideband frequency” measurement [7], where $\nu_{c}=\nu_{+}+\nu_{-}$ for an ideal Penning trap, $\nu_{-}$ and $\nu_{+}$ being the ion’s magnetron and reduced cyclotron radial eigenmotions respectively. This is done by obtaining the phases corresponding to the two radial motions acquired over periods of no excitation. The details of PI-ICR implementation at the CPT are described in Ref. [8]. One $\nu_{c}$ measurement cycle consists of two phase measurements: a reference phase ($\phi^{\text{ref}}$) and a final phase ($\phi^{\text{final}}$). Each of these phase measurements involves application of three RF excitations for each trap cycle: a $\nu_{-}$ dipole centering pulse, a $\nu_{+}$ dipole pulse to increase the radius of the corresponding motion and excite the ions to an orbit, and finally a $\nu_{c}$ quadrupole pulse to convert the ions’ fast $\nu_{+}$ motion to the slow $\nu_{-}$ motion for detection before they are ejected from the trap. This is repeated for multiple trap cycles to obtain enough ion-hits forming spots corresponding to the measured reference or final phase. The total trap time, durations of all three excitations, and timing structure of the two dipole excitations are kept the same for both the reference and the final phase measurements. The time of application of the $\nu_{c}$ pulse differentiates $\phi^{\text{ref}}$ and $\phi^{\text{final}}$. For the reference phase, the $\nu_{c}$ pulse is applied right after the end of the $\nu_{+}$ pulse, so the ions acquire $\nu_{-}$ phase in the time ($T$) between the conversion and the ejection pulse. As a result, the reference phase, in principle, is a $\nu_{-}$ phase. For the final phase, there is a delay ($t_{\text{acc}}$) between the two pulses, and the ions accumulate a combination of strongly mass-dependent $\nu_{+}$ phase over time $t_{\text{acc}}$, and $\nu_{-}$ phase over time $T-t_{\text{acc}}$. This means all beam species get resolved over a large-enough $t_{\text{acc}}$ for the final spots. The pulse schemes used for reference phase and final phase measurements are shown in Fig. 3(a), while a simulated spectrum of a reference and a final spot is shown in Fig. 3(b). It should be noted here that the conversion pulse preserves the accumulated phases of the eigenmotions, while flipping their signs [7, 19]. At the CPT, the ions’ eigenmotions take place in a clockwise direction, which materialize in a counter-clockwise phase accumulation from the reference to the final spot, as shown in Fig. 3(b). Each reference and final spot is clustered and fitted using techniques described in Ref. [20, 21] to obtain the corresponding spot-centroids and hence the acquired phases. A phase projection of the two spots is shown in Fig. 3(c). The phase difference $\phi_{\text{total}}=\phi^{\text{final}}-\phi^{\text{ref}}$ can then be determined, and from it, one can obtain $\nu_{c}$, without considering any systematic effects, as:
@@
Determination of $N$ forms an important part of precision measurement. From Eq. 3, a difference of $\Delta N$ turn would lead to a shift of $\Delta N/t_{\text{acc}}$ Hz in the determined $\nu_{c}$. For example, a difference of 1 turn for a measurement at $t_{\text{acc}}\sim 1000$ ms would lead to a relative shift of 1 ppm in the determined $\nu_{c}$. Therefore, conducting accurate measurements, especially at large $t_{\text{acc}}$, require prior knowledge of a precise guess $\nu_{c}$ to unequivocally determine $N$. For calibrant stable species, whose masses are precisely known, $N$ is obtained from a guess frequency, determined from the reported mass in AME2020 [15]. For target nuclides, a procedure is followed at the beginning of a measurement campaign to obtain the guess frequency. A series of final phase measurements are conducted, starting with a small $t_{\text{acc}}$ ($\sim$5 ms) and then gradually increasing it, to determine more precise and accurate values for $\nu_{c}$. The guess $\nu_{c}$ is updated with every such measurement. This is particularly necessary when measuring nuclides for which no prior experimental mass-data are available. For an impure beam, this method also allows for the identification of the $\nu_{c}$ frequencies of every species present in the beam, which is required for a phase correction (see Sec. 4.3.5). Once all beam species have been identified, a coarse $t_{\text{acc}}$ is selected ($\sim$100s of ms) such that all species in the beam are separated and clearly visible. Next a series of precise $t_{\text{acc}}$ values are selected, $\sim$100s of $\mu$s apart and spanning across roughly 1 - 1.5 $\nu_{-}$-periods, as required to correct for an initial $\nu_{-}$ motion (see Sec. 4.3.5). Additionally, each $t_{\text{acc}}$ is chosen such that $|\phi_{c}|\leq 10^{\circ}$, to account for the non-circular projection (see Sec. 4.3.1). The time $T$ is chosen such that during a final phase measurement, $\nu_{+}$ phase accumulation over $t_{\text{acc}}$ and complete conversion to $\nu_{-}$ motion via the $\nu_{c}$ pulse can take place. Typically, $T$ is set 5 - 10 ms longer than $t_{\text{acc}}$, and remains unchanged for all reference and final phase measurements for a given $\nu_{c}$ measurement.
@@ 4.3 Systematic Effects
A number of systematic effects has been observed during PI-ICR implementation at the CPT. Some of these have been outlined in Ref. [8], and continued to be refined. A few other effects like dependence on the angle of measurement of the target and calibrant species, characterization of the magnetic field and ion-ion interaction have been observed and studied since Ref. [8]. For completeness, a brief overview of all the effects is provided here.
@@ 4.3.4 Ion-ion interaction
At the CPT, the number of trapped ions is controlled by limiting the rate of incoming ions into the trap. The effect of ion-ion interaction was studied by conducting $\nu_{c}$ measurement of 133Cs+ ions from SIS using five datasets with varying ion rates between 0.5 pps to 10 pps on the PS-MCP with a $\sim$70detector efficiency, keeping all other conditions identical. The data was analysed using multiple ranges of ion-cut111ion-cut considers only the data involving a user-defined range for number of ion-hits detected simultaneously on the PS-MCP. condition. This is shown in Fig. 9, along with a linear fit with a slope of $-6(1)\times 10^{4}$ Hz///ions detected. During measurement campaigns, the ion rates are carefully maintained at $\sim$1-2 pps. For most species from CARIBU, this was not a challenge, as they were detected at the CPT with rates $\leq 1$ pps. For other species that are more abundantly produced including the stable ions from CARIBU (namely 84Kr+, 86Kr+ and C6H+${}_{6}^{+}$) or ions from SIS, this would be done by detuning one or two steerers upstream from the linear trap. Furthermore, during offline analysis, the ion-cut condition is typically set to $1-4$, and care is taken such that same number of detected ions per shot are considered for the target and the calibrant species, minimizing this effect to around 1 ppb in mass measurements.
@@ 4.3.5 Other systematic effects
where $k_{1}$ and $k_{2}$ are the amplitude and phase of the $\nu_{c}$ oscillation. The baseline $\overline{\nu}_{c}$ obtained from the fit gives the true cyclotron frequency. For more scarcely produced isotopes, 1 - 2 $\nu_{c}$ measurements are conducted, about 1///2 $\nu_{-}$ period apart, and then the model parameters are characterized using an abundant isobar ion to deduce $\overline{\nu}_{c}$ of the target species, while inflating the measured uncertainty. Finally, effects related to the difference in $A/q$ between the calibrant and measured nuclide were estimated to be $\leq$0.4(2) ppb$/\Delta$u from measurements of isotopes of well-known masses [25] from SIS and CARIBU, where $\Delta u=(A/q)-(A/q)^{\text{Cal}}$. Detailed discussions of the mass dependent shift along with the angular dependence from Sec. 4.3.1 are being prepared for a future publication.
@@ 4.3.6 Summary
Mass dependent shift Ratio corrected as 0.4(2) ppb$/\Delta$u.

### v5

@@ 1 Introduction
The Canadian Penning Trap (CPT) [4, 5] is located at Argonne National Laboratory (ANL) where it has conducted precision mass measurements of neutron-rich nuclides from the CAlifornium Rare Isotope Breeder Upgrade (CARIBU) [6] at the Argonne Tandem Linac Accelerator System (ATLAS) facility, using the Phase Imaging Ion-Cyclotron-Resonance (PI-ICR) [7, 8, 9, 10, 11, 12] technique. This paper discusses the PI-ICR operations, and some new systematic studies which have enabled the CPT to achieve ppb-level precision in mass measurements.
@@ 3.4 Penning trap
Electrodes Voltages [V]
Capture Trap Eject
Bottom correction tube --17.5 --0.9413 --0.9413
Bottom end cap --13.5 --3.0836 --3.0836
Bottom correction ring --8.9622 --8.9622 --8.9622
Ring --12.7206 --12.7206 --12.7206
Top correction ring --8.9622 --8.9622 --8.9622
Top end cap --3.0836 --3.0836 --13.5
Top correction tube --0.9413 --0.9413 --17.5
@@ 4.1 PI-ICR at the CPT
PI-ICR relies on determining a nuclide’s $\nu_{c}$ via a “sideband frequency” measurement [7], where $\nu_{c}=\nu_{+}+\nu_{-}$ for an ideal Penning trap, $\nu_{-}$ and $\nu_{+}$ being the ion’s magnetron and reduced cyclotron frequency respectively. This is done by obtaining the phases corresponding to the two radial motions acquired over periods of no excitation. The details of PI-ICR implementation at the CPT are described in Ref. [8]. One $\nu_{c}$ measurement cycle consists of two phase measurements: a reference phase ($\phi^{\text{ref}}$) and a final phase ($\phi^{\text{final}}$). Each of these phase measurements involves application of three RF excitations for each trap cycle: a $\nu_{-}$ dipole centering pulse, a $\nu_{+}$ dipole pulse to increase the radius of the corresponding motion and excite the ions to an orbit, and finally a $\nu_{c}$ quadrupole pulse to convert the ions’ fast $\nu_{+}$ motion to the slow $\nu_{-}$ motion for detection before they are ejected from the trap. This is repeated for multiple trap cycles to obtain enough ion-hits forming spots corresponding to the measured reference or final phase. The total trap time, durations of all three excitations, and timing structure of the two dipole excitations are kept the same for both the reference and the final phase measurements. The time of application of the $\nu_{c}$ pulse differentiates $\phi^{\text{ref}}$ and $\phi^{\text{final}}$. For the reference phase, the $\nu_{c}$ pulse is applied right after the end of the $\nu_{+}$ pulse, so the ions acquire $\nu_{-}$ phase in the time ($T$) between the conversion and the ejection pulse. As a result, the reference phase, in principle, is a $\nu_{-}$ phase. For the final phase, there is a delay ($t_{\text{acc}}$) between the two pulses, and the ions accumulate a combination of strongly mass-dependent $\nu_{+}$ phase over time $t_{\text{acc}}$, and $\nu_{-}$ phase over time $T-t_{\text{acc}}$. This means all beam species get resolved over a large-enough $t_{\text{acc}}$ for the final spots. The pulse schemes used for reference phase and final phase measurements are shown in Fig. 3(a), while a simulated spectrum of a reference and a final spot is shown in Fig. 3(b). It should be noted here that the conversion pulse preserves the accumulated phases of the eigenmotions, while flipping their signs [7, 23]. At the CPT, the ions’ eigenmotions take place in a clockwise direction, which materialize in a counter-clockwise phase accumulation from the reference to the final spot, as shown in Fig. 3(b). Each reference and final spot is clustered and fitted using techniques described in Ref. [24, 25] to obtain the corresponding spot-centroids and hence the acquired phases. A phase projection of the two spots is shown in Fig. 3(c). The phase difference $\phi_{\text{total}}=\phi^{\text{final}}-\phi^{\text{ref}}$ can then be determined, and from it, one can obtain $\nu_{c}$, without considering any systematic effects, as:
@@
The determination of $N$ is crucial for conducting accurate measurements. From Eq. 3, a difference of $\Delta N$ turn would lead to a shift of $\Delta N/t_{\text{acc}}$ Hz in the determined $\nu_{c}$. For example, a difference of 1 turn for a measurement at $t_{\text{acc}}\sim 1000$ ms would lead to a relative shift of about 1.5 ppm in the determined $\nu_{c}$ for 133Cs+. Therefore, conducting accurate measurements, especially at large $t_{\text{acc}}$, require prior knowledge of a precise guess $\nu_{c}$ to unequivocally determine $N$. For calibrant nuclides, whose masses are precisely known, $N$ is obtained from a guess frequency, determined from the reported mass in AME2020 [19]. For target nuclides, a procedure is followed at the beginning of a measurement campaign to obtain the guess frequency. A series of final phase measurements are conducted, starting with a small $t_{\text{acc}}$ ($\sim$5 ms) and then gradually increasing it, to determine more precise and accurate values for $\nu_{c}$. The guess $\nu_{c}$ is updated with every such measurement. This is particularly necessary when measuring nuclides for which no prior experimental mass-data are available. For an impure beam, this method also allows for the identification of the $\nu_{c}$ frequencies of every species present in the beam, which is required for a phase correction (see Sec. 4.3.5). Once all beam species have been identified, a coarse $t_{\text{acc}}$ is selected ($\sim$100s of ms) such that all species in the beam are separated and clearly visible. Next a series of precise $t_{\text{acc}}$ values are selected, $\sim$100s of $\mu$s apart and spanning across roughly 1 - 1.5 $\nu_{-}$-periods, as required to correct for an initial $\nu_{-}$ motion (see Sec. 4.3.5). Additionally, each $t_{\text{acc}}$ is chosen such that $|\phi_{c}|\leq 10^{\circ}$, to account for the non-circular projection (see Sec. 4.3.1). The time $T$ is chosen such that during a final phase measurement, $\nu_{+}$ phase accumulation over $t_{\text{acc}}$ and complete conversion to $\nu_{-}$ motion via the $\nu_{c}$ pulse can take place. Typically, $T$ is set 5 - 10 ms longer than $t_{\text{acc}}$, and remains unchanged for all reference and final phase measurements for a given $\nu_{c}$ measurement.
@@ 4.3 Systematic Effects
A number of systematic effects have been observed during PI-ICR implementation at the CPT. Some of these have been outlined in Ref. [8], and continued to be refined. A few other effects like dependence on the angle of measurement of the target and calibrant species, characterization of the magnetic field and ion-ion interaction have been observed and studied since Ref. [8]. For completeness, a brief overview of all the effects is provided here.
@@ 4.3.4 Ion-ion interaction
At the CPT, the number of trapped ions is controlled by limiting the rate of incoming ions into the trap. The effect of ion-ion interaction was studied by conducting $\nu_{c}$ measurement of 133Cs+ ions from SIS using five datasets with varying ion rates between 0.5 pps to 10 pps on the PS-MCP with a $\sim$70detector efficiency, keeping all other conditions identical. The data was analysed using multiple ranges of ion-cut111ion-cut considers only the data involving a user-defined range for number of ion-hits detected simultaneously on the PS-MCP. condition. This is shown in Fig. 9, along with a linear fit with a slope of $-6(1)\times 10^{-4}$ Hz///ions detected. During measurement campaigns, the ion rates are carefully maintained at $\sim$1-2 pps. For most species from CARIBU, this was not a challenge, as they were detected at the CPT with rates $\leq 1$ pps. For other species that are more abundantly produced including the stable ions from CARIBU (namely 84Kr+, 86Kr+ and C6H+${}_{6}^{+}$) or ions from SIS, this would be done by detuning one or two steerers upstream from the linear trap. Furthermore, during offline analysis, the ion-cut condition is typically set to $1-4$, and care is taken such that same number of detected ions per shot are considered for the target and the calibrant species, minimizing this effect to around 1 ppb in mass measurements.
@@ 4.3.5 Other systematic effects
where $k_{1}$ and $k_{2}$ are the amplitude and phase of the $\nu_{c}$ oscillation. The baseline $\overline{\nu}_{c}$ obtained from the fit gives the true cyclotron frequency. For more scarcely produced isotopes, 1 - 2 $\nu_{c}$ measurements are conducted, about 1///2 $\nu_{-}$ period apart, and then the model parameters are characterized using an abundant isobar ion to deduce $\overline{\nu}_{c}$ of the target species, while inflating the measured uncertainty. Finally, effects related to the difference in $A/q$ between the calibrant and measured nuclide were estimated to be $\leq$0.4(2) ppb$/\Delta$u from measurements of isotopes of well-known masses from SIS and CARIBU, where $\Delta u=(A/q)-(A/q)^{\text{Cal}}$. Detailed discussions of the mass dependent shift along with the angular dependence from Sec. 4.3.1 are being prepared for a future publication.
@@ 4.3.6 Summary
Mass dependent shift Ratio corrected as 0.4(2) ppb$/\Delta$u. The shift is also added in quadrature to the uncertainty to be conservative.